OpenStudy (anonymous):

Using the elimination method how to solve the value of x and y. -6x-y=29 / 8x+3y=-47

OpenStudy (cwrw238):

multiply the first equation by 3 to give -18x - 3y = 87 now add the first one to eliminate y: 8x + 3y = -47

OpenStudy (cwrw238):

this gives an equation in x which you can solve

OpenStudy (cwrw238):

then plug the value of x in the original first equation to find y

OpenStudy (anonymous):

first to eliminat x mul ist equation with coefficient of x in 2nd term 2nd equation with coefficient of x in 1st term now add terms or subtract to get value of x

OpenStudy (anonymous):

How about using the substitution and comparison method? using the same given example

OpenStudy (anonymous):

now put value of y in 1st or 2nd term to get valu of x

OpenStudy (anonymous):

For substitution method: Find the value of y from first and then put it into second to get x..

OpenStudy (anonymous):

first to eliminat x mul ist equation with coefficient of x in 2nd equation 2nd equation with coefficient of x in 1st euation now add or subtract 2 equations to get value of x now put value of y in 1st or 2nd equation to get value of x

OpenStudy (anonymous):

multiply the first equation by 3 on both sides -18x - 3y = 87 now add the first equation to this equation to eliminate y: -18x - 3y = 87 8x + 3y = -47 ------------ u try next steps

OpenStudy (anonymous):

then plug the value of x in the first or second equation to find y

OpenStudy (anonymous):

I'm a little confused on how to get the value of y.

OpenStudy (anonymous):

-18x - 3y = 87 8x + 3y = -47 ------------ -10x=40 x=-4 now put in 2nd equation 8(-4) + 3y = -47 3y=-47+24 3y=-23 y=-23/3